the total no of m for the electrons n=4 is ?................................................
 

Grade:9

1 Answers

Deepak Kumar
6 years ago
For n=4, first we have 4 values of l = 0,1,2,3. Now, for each value of l , we will have different values of m. So, for l =0, m=0. For l=1 , m= -1,0,1. For l = 2 , m = -2,-1,0,1,2. For l= 3 , m= -3,-2,-1,0,1,2,3. So counting each m values , total m values for n=4 :- 1+3+5+7 =16